3. Find the equation of tangent line to the curve y = x2 – Vx at the point (-1,0). 4. At a certain

Find the equation of tangent line to the curve y = x2 – \sqrt[3]{x} at the point (-1,0). 
Answer #1

SOLUTION :


y = x^2 - cube root(x) = x^2 - x^(1/3)


So, slope of the tangent =  dy/dx = 2x - 1/3 x^(- 2/3)


At point (- 1, 0) the slope = dy/dx at x = - 1,  is :


= 2(-1) - 1/3 (-1)^(-2/3)

= - 2 - 1/3 * 1/((-1)^(2/3))

= - 2 - 1/3 * 1/(-1)^2

= - 2 - 1/3 

= -7/3 


So, equation of the tangent at point (- 1, 0) is :


y = - 7/3 x + c


As point (- 1, 0) is both on the curve and the tangent :

=> 0 = - 7/3 (-1) + c

=> c = - 7/3

So,

Equation of the tangent at point (-1, 0) is :


y = - 7/3 x - 7/3  (ANSWER).


answered by: Tulsiram Garg